Shift report 20010817 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Aug 17 2001 - 07:57:25 EDT

  • Next message: Apache: "Shift report 20010817 08:00-16:00"

    ------------------------------------------------------------
    Date: Fri Aug 17 07:57:25 2001, Shift: 00:00-08:00
    Supervisor: Alexandru JIPA, with: Andrei MAKEEV
    Log book#7 pages 107 - 109
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beams up to 1.00 am. Run 4246 stopped. At 3.00 am the
    beams were cogged and stored. Run 4251 up to 6.50 am (new
    problems with the beams).
    2. Two data runs: 4246 (since 20.37) and 4251 (3.14 am -
    6.50 am). Other runs for ZDCs modifications tests. For the
    second experimental data run some problems at the HPSS 
    (solved around 7.00 am).
    3. No problems with detectors. Some problems with the two 
    cameras during the check (images suddenly lost and camera 
    is blocked). AM made some works for ZDCs.
     
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Aug 17 2001 - 07:58:03 EDT